From: Mon 8/11/08


  • UofC ATLAS meeting (3rd HEP)
  • discussion with Robert:
  • FDR meeting (phone EsNet: call 1-510-883-7860, code 888726#; esnet video, ad hoc 888726)
  • ATLAS shift (light)
  • Troubleshoot HU_ATLAS_Tier2
  • DQ2 tests (cont)
    • multiple streams, reduced DS (100 files)


  • Prepare shift report
  • ATLAS ADCoS (prodsys shifters) meeting (phone CERN, Xavier Espinal)
  • ATLAS TAG developers meeting (phone CERN, David Malon)
  • MWT2 meeting (RI405)
  • Test utility to delete datasest:
    • it deleted the files from the SE, the entries from LRC but the wrong dataset from the central catalog (the dataset was still there and 2 other datasets with similar name were deleted)
    • email to Charles and MWT2 list
  • ATLAS shift (almost light)
  • Troubleshoot HU_ATLAS_Tier2
  • DQ2 tests (cont)
    • multiple requests from same host
    • multiple clients, multiple servers



  • deleted ~20GB of work data from tier2-06


  • mounted tier2-06 disk on a WN for recovery attempts

-- MarcoMambelli - 13 Aug 2008
Topic revision: r3 - 15 Oct 2008, MarcoMambelli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback